40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jan 9 20:01:25 2019, Jon, Update, VAC, Second pumpdown completed Screenshot_from_2019-01-09_20-00-39.pngIMG_3088.jpg
    Reply  Thu Jan 10 10:23:46 2019, gautam, Update, VAC, overnight leak rate OvernightLeak.png
       Reply  Sat Feb 2 21:53:24 2019, Koji, Update, VAC, overnight leak rate Screen_Shot_2019-02-02_at_21.49.33.png
          Reply  Sun Feb 3 15:15:21 2019, gautam, Update, VAC, overnight leak rate Jan10_data.pngFeb1_data.png
             Reply  Sun Feb 3 20:52:34 2019, Koji, Update, VAC, overnight leak rate 
                Reply  Mon Feb 4 12:23:24 2019, gautam, Update, VAC, pumpdown 83 - leak tests PD83.png
                   Reply  Tue Feb 5 10:11:30 2019, gautam, Update, VAC, leak tests complete, pumpdown 83 resumed Annuli.pngMainVol.png
                      Reply  Tue Feb 5 19:30:14 2019, gautam, Update, VAC, Main volume at 20 uTorr PD83.png
                         Reply  Thu Feb 7 13:55:25 2019, gautam, Update, VAC, RGA turned on 
                            Reply  Thu Feb 7 19:28:46 2019, gautam, Update, VAC, IFO recovery 
Message ID: 14440     Entry time: Thu Feb 7 19:28:46 2019     In reply to: 14438
Author: gautam 
Type: Update 
Category: VAC 
Subject: IFO recovery 

[rana, gautam]

The full 1 W is again being sent into the IMC. We have left the PBS+HWP combo installed as Rana pointed out that it is good to have polarization control after the PMC but before the EOM. The G&H mirror setup used to route a pickoff of the post-EOM beam along the east edge of the PSL table to the AUX laser beat setup was deemed too flaky and has been bypassed. Centering on the steering mirror and subsequently the IMC REFL photodiode was done using an IR viewer - this technique allows one to geometrically center the beam on the steering mirror and PD, to the resolution of the eye, whereas the voltage maximization technique using the monitor port and an o'scope doesn't allow the former. Nominal IMC transmission of ~15,000 counts has been recovered, and the IMC REFL level is also around 0.12, consistent with the pre-vent levels.

ELOG V3.1.3-